# Interpretable LLM-based Table Question Answering

### v2

@@ Appendix A Explanation Methods for Table QA
In this section, we present visual explanations for Table QA models, which help bridge the gap between model behavior and human understanding. Each visualization provides insights into how the model interprets the input table, highlighting the key information used in its reasoning process.
@@
For our experiments, we use TabFact (Chen et al., 2020) and WikiTQ (Pasupat and Liang, 2015), running each method on the test sets of 2,024 and 4344 samples. POS, CoTable, and DATER use the same visualization representation as described in  Section 3.4.
@@
We present four different methods for explaining Table QA predictions: Text-to-SQL, DATER, CoTable, and POS (ours). Each method offers a unique set of information to visualizing explanations.
@@
  • Text-to-SQL (Rajkumar et al., 2022). This method translates the question into a single SQL command, which is then executed on the input table to produce an answer. Explanation Generation: We present the generated SQL command along with the question and the input table to show how the final answer is derived. Although the SQL itself can be clear to experts, it may require additional domain knowledge to interpret (see Figure 4 for an example).
@@
  • DATER (Ye et al., 2023). DATER decomposes a natural language query into subtasks and uses SQL to verify partial facts. Explanation Generation: We extract (1) the subtables selected by DATER from the input table, (2) the verified conditions in each subtask, and (3) attribution maps that highlight which table cells DATER judged as relevant. Additionally, we retrieve brief step descriptions (e.g., “Select Rows”) from DATER’s working logs, enabling users to track each subtask in the decomposition (see Figure 5).
@@
  • CoTable (Wang et al., 2024). CoTable processes queries by planning a sequence of abstract function calls (e.g., f_sort, f_select_row), each responsible for table manipulation. Explanation Generation: Similar to POS, we visualize each intermediate table along with attribution maps, but the transformations are represented by function names and their arguments (e.g., f_select_row(1)). Although this step-by-step approach can be informative, arguments are not justified and the final answer relies on LLM-driven “black-box” reasoning (see Figure 6).
  • POS (Ours). POS enforces atomic table transformations by generating small, single-condition SQL operations in a sequence, ensuring that each step’s impact on the table is transparent. Explanation Generation: We show each atomic step in natural language, convert it into an SQL query, and highlight the specific rows or cells selected in that step. This process yields direct insight into both why and how the table is transformed at each stage (see Figure 7).
@@
Comparison of DATER and CoTable. Empirically, we find no major difference in their overall “informativeness” (Section 4.1.3), although CoTable displays intermediate tables more explicitly. Both methods are still considered hybrid—they rely on partial SQL or function calls combined with LLM-driven reasoning. For tasks demanding deeper inspection of intermediate steps (e.g., verification complex filters), CoTable’s step-by-step interface may be more revealing than DATER’s subtable-based approach.
@@
In general, by generating method-specific explanations under a unified visualization scheme, we can fairly compare how each Table QA model interprets data and arrives at its final answer. This shared framework allows us to examine strengths, weaknesses, and interpretability trade-offs across different approaches in earlier sections.
A.1. Attribution Maps
@@
  • Adding the tracking index column: Before executing an SQL, we add a tracking index column to the current table. This column contains the original row indices from the initial table– Figure 3(a).
@@
  • Executing the SQL command: An SQL command is executed on the table with the tracking index column, producing a modified table– Figure 3(b).
@@
  • Identifying selected rows: After execution, we use the tracking index column to identify which rows have been selected or filtered by the SQL command– Figure 3(c).
@@
  • Identifying selected columns: We parse the SQL command to extract the columns involved in the operation– Figure 3(d) (see Appendix D).
@@
  • Visualizing an attribution map: The previous steps allow us to generate an attribution map for the initial table– Figure 3(e). The index column is also removed here.
@@
Figure 3. Generating attributions maps for POS. Column idx is added to track row attribution.
@@
Since both rows and columns can be attributed within an operation, POS offers a distinct advantage over previous works (Ye et al., 2023; Wang et al., 2024)—accurately attributing responsible cells for each transformation. For example, when an SQL command includes a condition that requires a cell to match a specific value or range (e.g., WHERE opponents = 0), we can determine which cells in the opponents column satisfy this condition and are thus responsible for the answer– Figure 3(e).
A.2. Explanations as Chains of Attribution Maps
@@
At each stage of the table transformation process, we visualize attribution maps on the intermediate table, emphasizing the data selected or filtered in the current operation. Rows and columns that contain relevant data for the operation are yellow-highlighted, while cells that match the specific condition in this step are green-highlighted. Using the information obtained from the plan execution and attribution maps, we combine the three components: (1) intermediate tables; (2) attribution maps; and (3) step description; to create an explanation. We present the explanation in a chain of attribution maps for Table QA models, helping users visually follow the sequence of transformations and understand how each feature contributes to the final answer.
@@
Figure 4. Text-to-SQL explanations provide only the SQL command, which is intuitive for domain experts.
@@
Figure 5. DATER explanations contain Sub-table Selection (S), contextual information (C), and highlights (H) that reveal which input features influence the prediction. In DATER, the cells used to construct the Sub-table Selection are yellow-highlighted. Additionally, the contextual information has been fact-checked using SQL commands against the input table.
@@
Figure 6. CoTable explanations present intermediate tables (T) and highlights (H), showing key steps in data transformation. In CoTable, intermediate tables and attributions (Nguyen et al., 2021b) are provided. Additionally, the steps are presented through function names and their arguments.
@@
Figure 7. POS (ours) explanation contains input Table T, input query Q, and prediction P, intermediate Table T, highlights H. The green-highlighted cells indicate where the information in the table matches the conditions specified in the natural language steps.
Appendix B Self-Explanation of Table QA Models
@@
An interesting question is whether a simple, post-hoc explanation method might achieve similar interpretability. To explore this, we adopt a widely used “self-explanation” baseline (Madsen et al., 2024; Chen et al., 2023b; Agarwal et al., 2024). Specifically, the LLM is first prompted to produce an answer and then asked to retrospectively explain how it arrived at that answer ( Figure 8). We compare this post-hoc approach with our proposed POS and other XAI baselines (Text-to-SQL, DATER, and CoTable) in Table 10, focusing on three XAI benchmarks in the TabFact dataset using GPT4o-mini.
@@
Table 10. Comparison of post-hoc self-explanation vs. other explanation methods. Lower is better for Preference; higher is better for Forward Simulation and Verification.
@@
Our main finding is that post-hoc self-explanation outperforms Text-to-SQL but remains significantly less effective than our proposed POS method. This may be attributed to its lack of faithfulness, as identified in previous studies (Madsen et al., 2024; Chen et al., 2023b), where the explanation often fails to align with the model’s true reasoning process. In contrast, POS ’s explanations are generated and executed at each step, inherently tying them to the actual transformations applied to the data. This design, as demonstrated in our earlier ablation and user studies, results in greater understanding and more actionable interpretability for downstream tasks like verification and forward simulation.
@@ 10 $9$-11-22 tennessee loss 6 13 7 - 3
To verify the statement "the wildcats kept the opposing team scoreless in four games," we need to examine the input table, which lists the results of the Kentucky Wildcats football games. The key information here is that the Wildcats are considered to have kept the opposing team scoreless if the column "opponents_points" shows a value of "0" for that game. …
@@
Figure 8. An example of a post-hoc self-explanation. The LLM first produces an answer (TRUE) and then justifies it by referencing the table entries that support the claim.
Appendix C Qualitative Examples for POS Explanations without Atomicity
@@
Below, we provide qualitative examples of POS explanations with and without atomicity in NL Planning. Removing atomicity from the plan steps can negatively impact interpretability, as the added complexity makes it harder for users to understand and trust the model’s reasoning process.
Figure 9. Upper: POS explanation with atomicity. Lower: POS explanation without atomicity.
@@
Figure 10. Two POS explanations without atomicity. The steps are compound and the attribution maps are non-trivial to comprehend.
Appendix D Extracting Columns from SQL Commands
@@
We design an algorithm to analyze SQL queries and identify the columns used within them.
Figure 11. Data-attribution tracking algorithm.
D.1. Algorithm Overview
@@ D.2. Implementation Details
D.3. An example of data-attribution tracking for Table QA
Here, we use the table transformation in Figure 23 as an example to illustrate our data-attribution tracking algorithm (Figure 11):

### v3

@@ Appendix A Explanation Methods for Table QA
In this section, we present visual explanations for Table QA models, which help bridge the gap between model behaviors and human understanding. Each visualization provides insights into how the model leverages the input table, highlighting the key information used in its reasoning process.
@@
For our experiments, we use TabFact (Chen et al., 2020) and WikiTQ (Pasupat & Liang, 2015), running each method on the test sets of 2,024 and 4344 samples. POS, CoTable, and DATER all use the same visualization format described in  Sec. 3.3.
@@
We use four different methods for explaining Table QA answers: Text-to-SQL, DATER, CoTable, and POS (ours). Each method offers a unique set of information as follows:
@@
  • Text-to-SQL (Rajkumar et al., 2022). This method translates the question into a single SQL command, which is then executed on the input table to produce an answer. Explanation Generation: We present the generated SQL command along with the question and the input table to show how the final answer is derived. Although the SQL itself can be clear to experts, it may require additional domain knowledge to interpret (see Fig. 4 for an example). Therefore, the we recruited Computer Science students (who have SQL expertise) for our human study.
@@
  • DATER (Ye et al., 2023). DATER solves a natural language question via extracting a relevant subtable and uses SQLs to verify partial facts (Fig. 5). Explanation Generation: We extract (1) the subtable selected by DATER from the input table, (2) the verified facts, and (3) attribution maps that highlight which table cells DATER considers relevant. Furthermore, we show step descriptions (e.g., “Select rows”) from DATER’s working logs, allowing users to track subtable extraction (see Fig. 5).
@@
  • CoTable (Wang et al., 2024). CoTable processes queries by planning a sequence of abstract function calls (e.g., f_sort, f_select_row, etc.), each responsible for table transformation. Explanation Generation: Similar to POS, we visualize each intermediate table along with attribution maps, but the transformations are represented by function names and their arguments (e.g., f_select_row(1)). Although this step-by-step approach can be informative, arguments are not well-justified and the final answer relies on LLM-driven “black-box” reasoning (see Fig. 6).
  • POS (Ours). POS decomposes a natural language question into a sequence of atomic reasoning steps, where each step is explicitly translated into an executable SQL command. Explanation Generation: For POS, we present the natural language description of each step—these are functionally equivalent to the underlying SQL commands, but far more accessible and readable to users. Users can inspect each transformation through the clear and concise natural language steps, supported by attribution maps highlighting the relevant table cells for every operation (see Fig. 7).
@@
Comparison of DATER and CoTable. Empirically, we find no major difference in their overall “informativeness” (Sec. 4.1.3), although CoTable displays intermediate tables more explicitly. Both methods are still considered hybrid—they rely on partial SQLs or function calls combined with LLM-driven reasoning. For tasks demanding deeper inspection of intermediate steps (e.g., verification complex filters), CoTable’s step-by-step interface may be more revealing than DATER’s subtable-based approach.
@@
In general, by generating method-specific explanations under a unified visualization format, we can fairly compare how each Table QA methods explain reasoning and arrives at its final answer. This shared framework allows us to examine strengths, weaknesses, and interpretability trade-offs across different approaches in earlier sections.
A.1 Attribution Maps
@@
  • Adding the tracking index column: Before executing an SQL, we add a tracking index column to the current table. This column contains the original row indices from the initial table— Fig. 3(a).
@@
  • Executing the SQL command: An SQL command is executed on the table with the tracking index column, producing a modified table— Fig. 3(b).
@@
  • Identifying selected rows: After execution, we use the tracking index column to identify which rows have been selected or filtered by the SQL command— Fig. 3(c).
@@
  • Identifying selected columns: We parse the SQL command to extract the columns involved in the operation— Fig. 3(d) (see Appendix E).
@@
  • Visualizing an attribution map: The prior information allow us to generate an attribution map for the intermediate tables— Fig. 3(e). The index column is also removed at this step.
@@
Figure 3: Generating attributions maps for POS. Column idx is added to track row attribution.
@@
Since both rows and columns can be attributed within an operation, POS offers a distinct advantage over previous works (Ye et al., 2023; Wang et al., 2024)—accurately attributing responsible cells for each transformation. For example, when an SQL command includes a condition that requires a cell to match a specific value or range (e.g., WHERE opponents = 0), we can determine which cells in the opponents column satisfy this condition and are thus responsible for the answer— Fig. 3(e).
A.2 Explanations as Chains of Attribution Maps
@@
At each step of the table transformation process, we visualize attribution maps on the input table for that step, highlighting the data selected or filtered in the current operation. Rows and columns that contain relevant data for the operation are yellow-highlighted, while cells that match the specific condition in this step are green-highlighted.
Using the information obtained from the plan execution and attribution maps, we combine the three components: (1) intermediate tables; (2) attribution maps; and (3) step description; to create an explanation. We present the explanation in a chain of attribution maps, helping users visually follow the sequence of transformations and understand how each table cell contributes to the final answer. We show representative examples of explanation methods in our study below.
@@
Figure 4: Text-to-SQL explanations provide only the SQL command, which is intuitive for SQL users.
@@
Figure 5: DATER explanations contain sub-table selection, contextual information (or verified facts), and highlights that reveal which input features influence the final answer.
@@
Figure 6: CoTable explanations present intermediate tables and highlights, showing key steps in data transformation. Additionally, the steps are presented through function names and their arguments.
@@
Figure 7: POS explanations contain intermediate tables and highlights. The green-highlighted cells indicate where the information in the table matches the conditions specified in the natural language steps.
A.3 Self-Explanation of Table QA Models
@@
An interesting question is whether a simple, post-hoc explanation method might achieve similar interpretability. To explore this, we adopt a widely used “self-explanation” baseline (Madsen et al., 2024; Chen et al., 2023b; Agarwal et al., 2024). Specifically, the LLM is first prompted to produce an answer and then asked to retrospectively explain how it arrived at that answer (see Fig. 8). We compare this post-hoc approach with our proposed POS and other XAI baselines (Text-to-SQL, DATER, and CoTable) in Tab. 10, focusing on three XAI benchmarks in the TabFact dataset using GPT-4o-mini.
@@
Table 10: Comparison of post-hoc self-explanation vs. other explanation methods. Lower is better for Preference; higher is better for Forward Simulation and Verification.
@@
Our main finding is that post-hoc self-explanation outperforms Text-to-SQL (always the worst) but remains significantly less effective than our proposed POS method. This may be attributed to its lack of faithfulness, as extensively identified in previous studies (Chen et al., 2023b; Madsen et al., 2024; Agarwal et al., 2024), where the explanation often fails to align with the model’s true reasoning process. In contrast, POS explanations are generated and executed at each step, inherently tying them to the actual transformations applied to the data. This design, as demonstrated in our earlier ablation and user studies, results in greater understanding and more actionable interpretability for downstream tasks like verification and forward simulation. We show an example of self-explanation below.
@@ 10 $9$-11-22 tennessee loss 6 13 7 - 3
To verify the statement "the wildcats kept the opposing team scoreless in four games", we need to examine the input table, which lists the results of the Kentucky Wildcats football games. The key information here is that the Wildcats are considered to have kept the opposing team scoreless if the column "opponents_points" shows a value of "0" for that game. …
@@
Figure 8: An example of a post-hoc self-explanation. The LLM first produces an answer (TRUE) and then justifies it by referencing the table entries that support the claim.
Appendix B Baselines and Related Works
LLM-only.
Program-only.
Program-based approaches generate explicit programs to perform table transformation and answer the question. Latent Program Algorithm (LPA) (Chen et al., 2020) frames TabFact verification as a program synthesis task, converting input queries into sequential operations (e.g., min, max, count, filter) executed via Python-Pandas. On the other hand, Text-to-SQL (Rajkumar et al., 2022) translates a natural language question directly into a single SQL command, which is then applied to the input table to generate the answer.
Hybrid.
Hybrid approaches combine the strengths of LLM reasoning and programs to perform Table QA and achieve state-of-the-art accuracy. DATER (Ye et al., 2023) uses an LLM to extract relevant sub-tables, while breaking queries into sub-queries and executing SQL commands to retrieve factual information. Binder (Cheng et al., 2023) takes a different approach by converting natural language questions into executable programs. It blends API calls with symbolic language interpreters like SQL or Python to address reasoning gaps that cannot be handled through offline methods alone. Lastly, CoTable (Wang et al., 2024) dynamically plans a sequence of predefined table operations–such as selecting rows or adding columns, allowing it to iteratively transform the table based on the intermediate information. Despite their differences, they all share a common strategy: they input the final simplified table along with the original question into an LLM to produce the final answer.
Program-aided language models vs. Plan-of-SQLs.
Although both PAL (Gao et al., 2023) and POS break questions into executable, program-based steps, they differ fundamentally in focus. PAL targets general math and symbolic problems by emitting a single Python script—comments and code intertwined—making it hard for non-programmers to trace each operation. In contrast, POS is built specifically for Table QA: it forces every reasoning step to be “atomic” (one condition and one variable at a time), translating each into a trivial SQL query. This guarantees that users can inspect and verify each intermediate result before moving on. Moreover, POS produces fine-grained attribution maps over the actual tables at every step, so users see exactly which rows and columns were involved—whereas PAL’s logic remains buried in Python code, opaque to those without coding expertise.
Appendix C More Ablation, Discussion, and Future Works
What is the impact of removing SQL execution from POS? We investigate this impact in Sec. 4.1.6. Removing SQL execution not only diminishes interpretability, but also affects accuracy on benchmark datasets. In this ablation, the LLM is tasked with directly transforming the table rather than executing SQL commands. While this leads to a negligible decrease in Table QA accuracy on TabFact, it causes a substantial drop on WikiTQ. This discrepancy suggests that relying solely on the black-box reasoning of LLMs for table transformations can severely impact model accuracy—likely due to LLM hallucinations or errors when handling complex tables (Chen, 2023; Wang et al., 2024). Moreover, bypassing SQL removes a layer of transparency, as the table transformations are no longer traceable, a reduction in interpretability that is further quantified by Tab. 5.
Have we evaluated explanation effectiveness in free-form Table QA? No; our work has not yet investigated the interpretability in free-form Table QA (e.g., FeTaQA), where the final answer is generated through a black-box operation of LLMs—making POS explanations no longer 100% faithful or comprehensive. A promising explanation approach is to develop hybrid explanations that blend POS-generated intermediate-step explanations with groundings (Hendricks et al., 2018) for the final answer. In this hybrid format, intermediate steps generated by POS provide a transparent view into the intermediate reasoning process while grounding algorithms are employed to anchor the final response to the intermediate tables.
What tabular questions are not SQL-decomposable? There are questions that cannot be broken down into Plans of SQLs (or sequential SQL queries). To our knowledge, they are:
  • Free-form narrative queries: e.g., “Summarize the basic information of the football clubs in Saint Petersburg”. They require surface-level composition or paraphrasing, not a relational result set.
  • World-knowledge or multi-table reasoning outside the given database: e.g., “Which of these universities is the oldest in the country”? Suppose that the table omits founding years. They require lookup beyond the input table; POS atomic SQL over the local table is, by construction, blind to external facts.
To automatically detect questions that are not suitable for POS, there are simple auto-detection strategies:
  • Rule-based heuristics: This is the most naive solution. We flag questions containing key verbs such as “summarize”, “describe”, or “compare” as likely free-form or analytic.
  • Execution-failure signals: Run the Step-to-SQL execution; if step converts to invalid SQL (Step-to-SQL raises error), flag as non-decomposable. This uses the same execution-validation hooks we have already described for error handling in Sec. 5.
  • LLM-based classification: We can prompt the NL planner to emit a Boolean “SQL-decomposable?” tag or fine-tune a lightweight classifier on a small labeled dataset of decomposable vs. non-decomposable questions.
SQL or Python; which is better for Table QA?
To see the difference, consider the query “top 3 departments by average salary among active employees over 40”. A Python solution might look like:
data = load_table("employees.csv")
step1 = [r for r in data if r["age"] > 40]
step2 = [r for r in step1 if r["status"] == "active"]
# custom code to group by department and compute averages
# then sort and slice the top 3
Using Python here has a few drawbacks:
  • Memory & speed: We must load the entire table into Python and loop over every row, which is slow and memory-intensive for large datasets.
  • Custom logic: Grouping, aggregating, and sorting require handwritten code (counters, loops, temporary structures) instead of built-in operations.
  • No indexing: Python loops cannot leverage database indexes or optimized query planners, leading to poor scalability.
SQL is more preferable for Table QA because:
  • Declarative clarity: We specify what we want (e.g., “Select rows where age > 40”), instead of using 3 steps sequentially: loop, filter > 40, and put to the list. This makes program more concise, self-documenting, and readable than using Python.
  • Engine optimizations: Modern database engines have extensively optimized query planning, and indexing; they can process large tables far faster and more memory-efficiently than Python loops [^3] https://airbyte.com/blog/sql-vs-python-data-analysis.
  • Consistency: SQL’s syntax and semantics are stable across databases (e.g., Oracle, MySQL, Microsoft SQL Server, and PostgreSQL), so the same query runs everywhere. Python script, by contrast, require custom connectors, custom layers, and may vary with library versions.
Ablation study on POS accuracy
In addition to the ablation study we conducted on interpretability (in Sec. 4.1.6), we run another ablation study on the accuracy of POS on both TabFact and WikiTQ to pinpoint modules which may not be necessary for either interpretability or performance.
Table 11: Accuracy (%) on TabFact and WikiTQ using GPT-4o-mini.
Variant TabFact (%) WikiTQ (%)
POS (Full) 82.70 59.32
– Atomic operations 78.11 (–4.59) 50.02 (–9.30)
– NL Planning 77.90 (–4.80) 48.27 (–11.05)
– SQL execution 79.84 (–2.86) 27.05 (–32.37)
We find that:
  • Replacing POS’s SQL transformations with end-to-end LLM calls causes the biggest accuracy loss—especially on WikiTQ (59.32 → 27.05%, –32.37 pp) and a smaller drop on TabFact (82.70 → 79.84%, –2.86 pp).
  • Removing either the natural-language planning or atomic-step enforcement each costs approximately 4–5 pp on TabFact and 9–11 pp on WikiTQ.
While all three components are important, robust SQL execution has the largest impact on accuracy, followed by step-by-step NL planning, and then atomic decomposition.
Appendix D Qualitative Examples for POS Explanations without Atomicity
@@
In this section, we provide qualitative examples of POS explanations with and without atomicity enforcement in NL Planning. Removing atomicity from the plan steps can negatively impact interpretability, as the added complexity makes it harder for users to understand the model’s reasoning process.
Figure 9: Upper: POS explanation with atomicity. Lower: POS explanation without atomicity.
@@
Figure 10: Two POS explanations without atomicity. The steps are compound and the attribution maps are non-trivial to comprehend.
Appendix E Extracting Columns from SQL Commands
@@
In this section, we detail the algorithm to analyze SQL queries and identify the columns used within them.
Figure 11: Data-attribution tracking algorithm for POS.
E.1 Algorithm Overview
@@ E.2 Implementation Details
E.3 An example of data-attribution tracking for Table QA
Here, we use the table transformation in Fig. 23 as an example to illustrate our data-attribution tracking algorithm (Fig. 11):